An automobile dealer sells 2 models of cars, A and B. Model A can be purchased in 7 different colors and 4 different engine sizes. Model B comes in 8 colors and 3 engines sizes. How many cars must the dealer order to have one car of each model in each color and engine size?
Answer by ikleyn(52788)

.

7*4 = 28 cars of model A and 8*3 = 24 cars of model B.


In total, 28 + 24 = 52 cars.    

Solved.